Consecutive Integers-Find the three consecutive integers whose sum is 84. I know the formula is n+(n+1)+(n+2)=84 and the answer should be 27,28,29 but I don't know how to show it.
Please help!

Answer by scott8148(5891)

3n + 3 = 84

3n = 81

n = 27

n + 1 = 28

n + 2 = 29